# DreamRunner: Fine-Grained Compositional Story-to-Video Generation with Retrieval-Augmented Motion Adaptation

### v3

@@ A.3 Region-Based LoRA Injection
Low-Rank Adaptation (LoRA) [22] was introduced to efficiently adapt large pre-trained language models to various downstream tasks. More recently, LoRA has been extended to text-to-image and text-to-video generation, enabling lightweight subject and motion customization [50, 75]. Instead of updating the full weight matrix $W$, LoRA applies a low-rank decomposition, modifying it as:
@@
Similarly, the witch’s LoRA only modifies the witch-related region ($Mask_{\text{witch}}\cdot x$), preserving consistency with traditional LoRA-based adaptation for the witch. For other regions unrelated to both the witch and the cat, the output remains unchanged, as the mask ensures that no LoRA modifications are applied, making it equivalent to the output without any LoRA injection. For the Motion LoRAs in main paper’s Fig. 1, we follow the similar style to inject them locally:
@@
Appendix B Additional Ablations
@@ B.1 Ablations of RAG pipeline
We provide additional experiments in Section 3.2 to demonstrate the effectiveness of our data processing approach within the retrieval pipeline. Table 5 primarily evaluates the impact of the maximal number of retrieved videos and the use of CLIP and ViCLIP for filtering. For efficiency, we evaluate a subset of eight motions selected from the full pool of 64 motions. The results indicate that, compared to the CogVideoX zero-shot baseline (Row 1), retrieving videos without any filtering (Row 2) improves performance, even though BM-25 [49] retrieval introduces some noise. This highlights the importance of retrieval itself. Furthermore, adding CLIP and ViCLIP as filters (Row 4) further enhances performance, showcasing the benefit of using semantically aligned videos for improved motion learning. Additionally, retrieving a sufficient number of videos is critical, as evidenced by Row 3, where limiting retrieval to a maximum of three videos results in poorer performance compared to retrieving 20 videos (Row 4).
@@ Appendix C Evaluation Metrics
  • Similarity to Reference Images: We assess the alignment between the generated videos and reference images using Image-Image CLIP [19] and DINO [42] scores. The evaluation is conducted by averaging the CLIP/DINO similarity scores between each frame of the generated video and the reference images, with CLIP-L14 (336px) [19] and DINOv2-B16 [42] as image encoders.
  • Similarity to Full Narration per Scene: We measure the alignment between the full narration and the generated videos for each scene using Image-Text CLIP [19] and Video-Text ViCLIP [58] scores. For the CLIP score, we uniformly sample eight frames from the single-scene video and compute the average score between each frame and the full narration. For the ViCLIP score, we directly use the alignment score between the video and the narration. Per-scene scores are averaged to obtain the overall score for the multi-scene storytelling video.
  • Fine-Grained Text Alignment per Scene: We also assess fine-grained alignment to textual descriptions using Image-Text CLIP [19] and Video-Text ViCLIP [58] scores. In our story, each narration contains two motions, which we decouple into two consecutive single-motion descriptions using an LLM [41]. For each generated single-scene video, we divide it into two segments at the temporal midpoint. We then compute the CLIP/ViCLIP scores between the first segment and the first description and between the second segment and the second description. The average of these two scores provides the per-scene score, and the per-scene scores are further averaged to compute the overall score for the storytelling video.
@@
  • Motion Smoothness: Following [23], we utilize motion priors from a video frame interpolation model [34] to quantify the smoothness of motion in generated videos (see Supplementary for details), specifically, given one video, we drop the odd-number frames and then use the interpolation model to infer the dropped frames. Then we compute the Mean Absolute Error between the estimated frames and the dropped frames.
  • Aesthetic Quality: We evaluate the artistic and visual appeal of each generated video frame using the LAION aesthetic predictor [51] frame-by-frame. This metric reflects aesthetic aspects such as composition, color richness and harmony, photorealism, naturalness, and overall artistic quality.
  • Imaging Quality: Imaging quality measures distortions such as over-exposure, noise, and blur in generated frames. We evaluate this using the MUSIQ [27] image quality predictor frame-by-frame, which is trained on the SPAQ dataset [12].
Appendix D Compositional T2V Examples
In this section, we present qualitative examples demonstrating the capabilities of DreamRunner in compositional text-to-video generation. DreamRunner effectively generates videos with accurate action binding to different characters, consistent attributes across objects, dynamic attribute changes, motion control, object interactions, and appropriate spatial relationships between objects.
@@ Appendix D Compositional T2V Examples
Appendix E Characters.
We provide character examples in Figure 11, where the first four are generated using FLUX [4], and the others are collected from existing customization datasets (CustomConcept101 [30] and Dreambooth [50]).
Appendix F Single-Character Examples
@@
Appendix G Multi-Character Examples
@@
Appendix H LLM Prompts
We provide detailed LLM prompts for both high-level plans and fine-grained plans in Listing LABEL:lst:hlplan and Listing LABEL:lst:fgplan. respectively. For high-level plans, we use a simple in-context example with instructions, while fine-grained plans require reasoning before generating the output. Example outputs are shown in Listings LABEL:lst:hleg and LABEL:lst:fgeg. For multi-character scenarios, we modify character-related words and examples and limit required motions to a maximum of four.
@@ Appendix H LLM Prompts
Listing 2: The example of the output from the high-level planning prompt, LABEL:lst:hlplan. We highlight scene, motions, narrations using same colors as Section 3.1.
@@
Listing 4: The example output of fine-grained planning prompt LABEL:lst:fgplan. We highlight Background, [entity, motion, caption] and layout [left, top, right, bottom] using same colors as Section 3.1.

### v4

@@ A.3 Region-Based LoRA Injection
Low-Rank Adaptation (LoRA) (hu2021lora) was introduced to efficiently adapt large pre-trained language models to various downstream tasks. More recently, LoRA has been extended to text-to-image and text-to-video generation, enabling lightweight subject and motion customization (ruiz2023dreambooth; zhao2023motiondirector). Instead of updating the full weight matrix $W$, LoRA applies a low-rank decomposition, modifying it as:
@@
Similarly, the witch’s LoRA only modifies the witch-related region ($Mask_{\text{witch}}\cdot x$), preserving consistency with traditional LoRA-based adaptation for the witch. For other regions unrelated to both the witch and the cat, the output remains unchanged, as the mask ensures that no LoRA modifications are applied, making it equivalent to the output without any LoRA injection. For the Motion LoRAs in main paper’s Fig. 1,, we follow the similar style to inject them locally:
@@
Appendix B Additional Comparisons
@@ B.1 Ablations of RAG pipeline
We provide additional experiments in Section 3.2 to demonstrate the effectiveness of our data processing approach within the retrieval pipeline. Table 5 primarily evaluates the impact of the maximal number of retrieved videos and the use of CLIP and ViCLIP for filtering. For efficiency, we evaluate a subset of eight motions selected from the full pool of 64 motions. The results indicate that, compared to the CogVideoX zero-shot baseline (Row 1), retrieving videos without any filtering (Row 2) improves performance, even though BM-25 (robertson2009probabilistic) retrieval introduces some noise. This highlights the importance of retrieval itself. Furthermore, adding CLIP and ViCLIP as filters (Row 4) further enhances performance, showcasing the benefit of using semantically aligned videos for improved motion learning. Additionally, retrieving a sufficient number of videos is critical, as evidenced by Row 3, where limiting retrieval to a maximum of three videos results in poorer performance compared to retrieving 20 videos (Row 4).
@@ B.2 Ablations of Odd-Even Layer Separation for Motion Prior Learning
B.3 Computational Cost Comparison
Since our method is training-free and only involves test-time fine-tuning (TTF), we report the optimization time solely for each method for reference. Specifically, our methods’ TTF takes approximately 0.2 GPU hours per motion, totaling 3–4 GPU hours for a full story with 15–20 motions.
In comparison, training-intensive baselines, VLogger and VideoDirectorGPT, require around 6K and 400 GPU hours, respectively. Notably, our method still outperforms these baselines even without RAG and TTF (Table 3), with only spatial-temporal regional attention involved.
@@ Appendix C Evaluation Metrics
  • Similarity to Reference Images: We assess the alignment between the generated videos and reference images using Image-Image CLIP (hessel2021clipscore) and DINO (oquab2023dinov2) scores. The evaluation is conducted by averaging the CLIP/DINO similarity scores between each frame of the generated video and the reference images, with CLIP-L14 (336px) (hessel2021clipscore) and DINOv2-B16 (oquab2023dinov2) as image encoders.
  • Similarity to Full Narration per Scene: We measure the alignment between the full narration and the generated videos for each scene using Image-Text CLIP (hessel2021clipscore) and Video-Text ViCLIP (wang2023internvid) scores. For the CLIP score, we uniformly sample eight frames from the single-scene video and compute the average score between each frame and the full narration. For the ViCLIP score, we directly use the alignment score between the video and the narration. Per-scene scores are averaged to obtain the overall score for the multi-scene storytelling video.
  • Fine-Grained Text Alignment per Scene: We also assess fine-grained alignment to textual descriptions using Image-Text CLIP (hessel2021clipscore) and Video-Text ViCLIP (wang2023internvid) scores. In our story, each narration contains two motions, which we decouple into two consecutive single-motion descriptions using an LLM (openai_gpt4o). For each generated single-scene video, we divide it into two segments at the temporal midpoint. We then compute the CLIP/ViCLIP scores between the first segment and the first description and between the second segment and the second description. The average of these two scores provides the per-scene score, and the per-scene scores are further averaged to compute the overall score for the storytelling video.
@@
  • Motion Smoothness: Following (huang2023vbench), we utilize motion priors from a video frame interpolation model (li2023amt) to quantify the smoothness of motion in generated videos (see Supplementary for details), specifically, given one video, we drop the odd-number frames and then use the interpolation model to infer the dropped frames. Then we compute the Mean Absolute Error between the estimated frames and the dropped frames.
  • Aesthetic Quality: We evaluate the artistic and visual appeal of each generated video frame using the LAION aesthetic predictor (laion_aesthetic_predictor) frame-by-frame. This metric reflects aesthetic aspects such as composition, color richness and harmony, photorealism, naturalness, and overall artistic quality.
  • Imaging Quality: Imaging quality measures distortions such as over-exposure, noise, and blur in generated frames. We evaluate this using the MUSIQ (ke2021musiq) image quality predictor frame-by-frame, which is trained on the SPAQ dataset (fang2020cvpr).
Appendix D Visual Quality Evaluation for SVG
We evaluate the generated videos using a comprehensive set of visual quality and consistency metrics, including aesthetics, imaging quality, motion smoothness, temporal flickering, subject consistency, and background consistency, as shown in Table 7 and Table 8. Our method achieves the best overall performance across all metrics, demonstrating its effectiveness in producing high-quality and consistent videos. Moreover, the ablation results indicate that introducing the proposed RAG and SR3AI modules does not harm the visual quality, as the full model maintains or improves performance compared to its variants.
Figure 5: Generated multi-character videos with slightly overlapping regions.
  Method Aesthetics Imaging Smoothness Temp. Flickering Subject Consistency BG Consistency Overall
VideoDirectorGPT (lin2023videodirectorgpt) 42.3 60.3 94.3 91.5 75.6 87.8 75.3
VLogger (zhuang2024vlogger) 43.4 61.2 96.2 95.7 84.1 91.0 78.6
DreamRunner (Ours) 55.4 (+27.65%) 62.1 (+1.47%) 98.1 (+1.98%) 96.2 (+0.52%) 91.2 (+8.44%) 92.3 (+1.43%) 82.6 (+5.03%)
Table 7: Evaluation of visual quality and consistency. Metrics include aesthetics, imaging, motion smoothness, temporal stability, subject/background consistency, and overall quality. DreamRunner Achieves the best quality compared with others.
  RAG SR3AI Aesthetics Imaging Smoothness Temp. Flickering Subject Consistency BG Consistency Overall
$\times$ $\times$ 54.3 61.3 94.3 95.9 91.0 92.2 81.50
$\times$ 55.4 61.9 98.0 96.3 90.9 92.3 82.47
$\times$ 55.6 61.9 98.1 96.0 91.1 92.0 82.45
55.4 62.1 98.1 96.2 91.2 92.3 82.55
Table 8: Ablation on quality dimensions. Overall is the mean of all six metrics. Adding RAG and SR3AI doesn’t harm the overall visual quality of the backbone model. The last row is our full method DreamRunner.
Appendix E Compositional T2V Examples
In this section, we present qualitative examples demonstrating the capabilities of DreamRunner in compositional text-to-video generation. DreamRunner effectively generates videos with accurate action binding to different characters, consistent attributes across objects, dynamic attribute changes, motion control, object interactions, and appropriate spatial relationships between objects. Note that in the overlapped regions, the caption is a merge of all involved captions. For cleaner visualization, we don’t show the merged caption in the layout plans.
@@ Appendix E Compositional T2V Examples
Appendix F Characters
We provide character examples in Figure 12, where the first four are generated using FLUX (flux), and the others are collected from existing customization datasets (CustomConcept101 (kumari2023multi) and Dreambooth (ruiz2023dreambooth)).
Appendix G Single-Character Examples
@@
Appendix H Multi-Character Examples
@@
Appendix I LLM Prompts
We provide detailed LLM prompts for both high-level plans and fine-grained plans in Listing 1 and Listing 3. respectively. For high-level plans, we use a simple in-context example with instructions, while fine-grained plans require reasoning before generating the output. Example outputs are shown in Listings 2 and 4. We also provide our example input output for merging captions in overlapped regions in Listing  5. For multi-character scenarios, we adapt character-specific words and examples, limit required motions to a maximum of four, and enforce non-overlapping spatial regions for each character.
Appendix J Multi-Character Video Generation with Overlapping Regions
We show an example of DreamRunner handling overlapped regions in a multi-character setting (see Fig. 5). In this case, the bounding boxes of the cat and the witch slightly overlap, leading to visual overlap between the characters themselves. This demonstrates DreamRunner’s ability to manage such interactions.
Appendix K Limitation and Future Work
Limitations. DreamRunner build s on a diffusion-based video generation framework and trains lightweight adapters on frozen backbones. As such, its performance and output quality are inherently constrained by the capabilities of the underlying backbone models. If the backbone struggles with rare compositions, complex motions, or fine-grained details, DreamRunner may inherit these weaknesses.
Future work. Future improvements may come from adopting stronger and more generalizable backbone models to enhance the overall visual quality. Additionally, advanced LoRA-merging techniques such as CLoRA (meral2024clora) can be explored to better coordinate multiple character streams and improve quality in overlapping regions. More broadly, developing architectures with stronger compositional reasoning and multi-entity awareness remains an important direction.
@@ Appendix I LLM Prompts
Listing 2: The example of the output from the high-level planning prompt, from Listing 1. We highlight scene, motions, narrations using same colors as Section 3.1.
@@
Listing 4: The example output of fine-grained planning prompt 3. We highlight Background, [entity, motion, caption] and layout [left, top, right, bottom] using same colors as Section 3.1.
@@
Listing 5: The example input output for getting merged prompt for overlapping regions. The model input is in black while the output is in blue
Given multiple captions describing the same scene, merge them into a single, coherent caption that includes all key actions or entities, avoids redundancy, and maintains natural fluency.
[Captions]
1. A man is throwing a frisbee in the park.
2. A dog is running toward the frisbee with its tongue out.
[Merged Caption]
A man throws a frisbee in the park while a dog runs toward it with its tongue out.